OpenStudy (anonymous):

√8y^5 x √40y^2 can you help me simplify the product algebra 2

OpenStudy (anonymous):

√8y^5 x √40y^2 um...

OpenStudy (anonymous):

the "x" is a multiplication sign

OpenStudy (anonymous):

8y=5x+8 i think if not...

OpenStudy (anonymous):

you know the √ is a radical right?

OpenStudy (anonymous):

Do you mean: \[\sqrt{8y^5}\times \sqrt{40y^2}\]?

OpenStudy (anonymous):

Yes, thank you

OpenStudy (anonymous):

Use this rule for radicals: \[\sqrt{a}\times \sqrt{b}=\sqrt{ab}\]